# uncompress

### v2

@@ 3.8. Toroidal varieties and logarithmic geometry
As there is no other applications of logarithmic geometry in this paper except in the proof of Theorem 3.11, we refer the readers to [Kat89, Kat94, GR18, Ogu18] for details about logarithmic geometry. For more details regarding the relation between toroidal geometry and logarithmic geometry, see [Qu25].
All the log-structures in this paper are defined on Zariski sites. A log-scheme $(X,\mathcal{M})$ consists of a scheme $X$ and a log-structure $\alpha\colon\mathcal{M}\to\mathcal{O}_{X}$; see [Kat94, §1]. A log-variety is a log-scheme whose underlying scheme is a variety. To avoid confusions, when saying logarithmically smooth log-morphisms, we mean logarithmically smooth morphisms in the category of log-schemes defined in the sense of [Kat89, Kat94, Ogu18]. However, when saying a log smooth morphism, we mean a log smooth morphism from a couple as in §2.8. The applications of these terminologies should be clear from the context.
Let $X$ be a variety and $\iota\colon U\hookrightarrow X$ an inclusion of an open subset. Denote by $\mathcal{M}_{U/X}$ the sheaf of monoids $\mathcal{O}_{X}\cap\iota_{*}\mathcal{O}_{X}^{*}$. We call $\mathcal{M}_{U/X}$ the compactifying log-structure associated to the embedding $(U\subset X)$; cf. [Ogu18, §III.1.6]. The log-variety $(X,\mathcal{M}_{U/X})$ is logarithmically regular (or log-regular to abbreviate) if and only if $(U\subset X)$ is a strict toroidal embedding; cf. [Kat94, (1.7)] and [Qu25].
Let $(U_{X}\subset X),(U_{Y}\subset Y)$ be strict toroidal embeddings. Let $(U_{X}\subset X)\to(U_{Y}\subset Y)$ be a morphism of embeddings, that is, a morphism $f\colon X\to Y$ of varieties such that $f(U_{X})\subseteq U_{Y}$. Then a morphism of embeddings $(U_{X}\subset X)\to(U_{Y}\subset Y)$ is a dominant toroidal morphism if and only if the induced log-morphism $(X,\mathcal{M}_{U_{X}/X})\to(Y,\mathcal{M}_{U_{Y}/Y})$ of log-varieties is logarithmically smooth in the category of log-schemes; see [Den13, Qu25].
@@ 3.9. Relative toroidalisation
First we recall a theorem first proved by Abramovich and Karu in [AK00], which says that any projective, surjective morphism of varieties can be made logarithmically smooth after birational modifications. This result is generalised to varieties over non-closed fields of characteristic zero in [ADK13].
Theorem 3.10 ([ADK13, Theorem 1.1]).
Let $f\colon X\to B$ be a projective, surjective morphism of varieties. Let $Z\subset X$ be a proper closed subscheme. Then there exists a commutative diagram
@@
We note that the method proving this theorem in [AK00, ADK13] is non-canonical and even a smooth generic fibre of $X\to B$ can be modified by it. In [ATW20], it is shown that there is a relatively canonical toroidalisation method that keeps the toroidal locus of $f$ unchanged; for more details, we refer the readers to [ATW20, §1]. In the next result, we show that a normalised base change of a dominant toroidal morphism is also toroidal. For more details and generalisation of this result to log-structures on small étale sites, we refer the readers to [Qu25].
@@
$f\colon(U_{Y}\subset Y)\to(U_{Z}\subset Z)\,\text{ and }\,g\colon(U_{X}\subset X)\to(U_{Z}\subset Z)$
@@
$f\colon(Y,\mathcal{M}_{U_{Y}/Y})\to(Z,\mathcal{M}_{U_{Z}/Z})\,\text{ and }\,g\colon(X,\mathcal{M}_{U_{X}/X})\to(Z,\mathcal{M}_{U_{Z}/Z})$
@@
$\mathbf{\Omega}\coloneqq(\Omega,\mathcal{M}_{\Omega})\,\text{ and }\,\mathbf{\Omega}^{\operatorname{fs}}\coloneqq(\Omega^{\operatorname{fs}},\mathcal{M}_{\Omega^{\operatorname{fs}}})$
the fibre product of $(Y,\mathcal{M}_{U_{Y}/Y})$ and $(X,\mathcal{M}_{U_{X}/X})$ along $(Z,\mathcal{M}_{U_{Z}/Z})$ in the category of log-schemes and fs log-schemes respectively. By [Ogu18, III.2.1.2], the underlying scheme $\Omega$ is isomorphic to $Y\times_{Z}X$, and $\mathcal{M}_{\Omega}$ is a coherent log-structure on $\Omega$.
By [GR18, Remark 12.2.36 (ii)], the morphism of schemes $\Omega^{\operatorname{fs}}\to\Omega$ is a finite surjective morphism. By [GR18, Proposition 12.3.34] and the proof of [GR18, Proposition 12.2.35],
@@
is log-étale. As logarithmic smoothness is stable under composition and base change (see [GR18, Proposition 12.3.24]), the induced log-morphism of log-schemes
@@
is logarithmically smooth. Then by [Kat94, (8.2)], the log-scheme $\mathbf{\Omega}^{\operatorname{fs}}$ is log-regular. By [Kat94, (4.1)], $\Omega^{\operatorname{fs}}$ has normal irreducible components, hence $\Omega^{\operatorname{fs}}$ is a disjoint union of normal varieties. Thus, every irreducible component of $\Omega^{\operatorname{fs}}$ is a strict toroidal variety.
Let $U$ be the fibre product $U_{Y}\times_{U_{Z}}U_{X}$. By [Qu25, Lemma 6.2], $U$ is nonempty. Let $(U,\mathcal{M}_{U})$ be the log-scheme endowed with trivial log-structure. It is evident that $\mathcal{M}_{U}$ is the restriction of $\mathcal{M}_{\Omega}$ to $U$. Denote by $\Omega^{\operatorname{fs},*}$ the maximal open subset of $\Omega^{\operatorname{fs}}$ on which $\mathcal{M}_{\Omega^{\operatorname{fs}}}$ is trivial. Then by the construction in [GR18, Proposition 12.2.35], $U$ is also an open subscheme of $\Omega^{\operatorname{fs}}$ such that $\mathcal{M}_{U}=\mathcal{M}_{\Omega^{\operatorname{fs}}}|_{U}$, whence $U\subseteq\Omega^{\operatorname{fs},*}$. Moreover, by [GR18, Corollary 12.3.27], the scheme $U$ is nonsingular in the usual sense.
Step 2. Let $W^{\prime}$ be an irreducible component of $Y\times_{Z}X$ that dominates $Y$. By [Qu25, Lemma 6.2], the inverse image of $U_{X}$ to $W^{\prime}$ is a dense open subset of $W^{\prime}$. Let $W$ be the normalisation of $W^{\prime}$, and let $p\colon W\to Y$ and $q\colon W\to X$ be the projections. Then $p^{-1}(U_{Y})\cap q^{-1}(U_{X})$ is a nonempty dense open subset of $W$. Since $U$ is a nonsingular scheme, there is a unique irreducible component $U_{W}$ of $U$ contained in $W^{\prime}$. Then the normalisation $W\to W^{\prime}$ is an isomorphism over $U_{W}$. Thus, viewed as an open subset of $W$, $U_{W}=p^{-1}(U_{Y})\cap q^{-1}(U_{X})$. Let $W^{\prime\prime}$ be the irreducible component of $\Omega^{\operatorname{fs}}$ that contains $U_{W}$. As $U_{W}$ embeds into both $W^{\prime}$ and $W^{\prime\prime}$, the finite morphism $W^{\prime\prime}\to W^{\prime}$ is also birational. Since $\Omega^{\operatorname{fs}}$ has normal irreducible components, we see that $W^{\prime\prime}\cong W$.
Denote by $W^{*}\subseteq W$ the restriction of $\Omega^{\operatorname{fs},*}$ to $W$. By [GR18, Remark 12.2.8], $W^{*}$ is mapped into $U_{Y},U_{X}$ via $p\colon W\to Y,q\colon W\to X$. Then
@@
    [Base change functoriality] There is a non-empty open subset $U_{\mathfrak{B}}\subset\mathfrak{B}\setminus\mathfrak{Z}_{\mathfrak{B}}$ satisfying the following property. Let $i\colon C\to\mathfrak{B}$ be a morphism from a smooth curve. Assume that the image $i(C)$ is not entirely contained in the closed subset $\mathfrak{B}\setminus U_{\mathfrak{B}}$. Then $\mathfrak{X}\times_{\mathfrak{B}}C$ admits a unique irreducible component dominating $C$. Denote by $Y$ the reduction of this irreducible component of $\mathfrak{X}\times_{\mathfrak{B}}C$ with projection morphism $f\colon Y\to C$. Consider the diagram arising from normalised base change as follows:
@@
    where $C^{\prime}$ (respectively, $Y^{\prime}$) is the normalisation of the main component of $\mathfrak{B}^{\prime}\times_{\mathfrak{B}}C$ (respectively, of $\mathfrak{X}^{\prime}\times_{\mathfrak{B}^{\prime}}C^{\prime}$), and $U_{C^{\prime}}$ (respectively, $U_{Y^{\prime}}$) is the pullback of $U_{\mathfrak{B}^{\prime}}$ to $C^{\prime}$ (respectively, of $U_{\mathfrak{X}^{\prime}}$ to $Y^{\prime}$). Then $f^{\prime}$ is also a toroidal morphism of toroidal embeddings and all the properties in (i) are satisfied. In particular, $m_{C}$ and $m_{Y}$ are projective birational morphisms, and the inverse image of $\mathfrak{Z}_{\mathfrak{B}}$ in $C^{\prime}$ (respectively, of $\mathfrak{Z}_{\mathfrak{X}}$ in $Y^{\prime}$) is contained in $C^{\prime}\setminus U_{C^{\prime}}$ (respectively, in $Y^{\prime}\setminus U_{Y^{\prime}}$).
@@
Step 1. To include the closed subset $\mathfrak{Z}_{\mathfrak{X}}$ into the toroidal structure, we first modify $\Phi\colon\mathfrak{X}\to\mathfrak{B}$ by Theorem 3.10 as in the following diagram
@@
where $m_{\mathfrak{X}^{\circ}}$ and $m_{\mathfrak{B}^{\circ}}$ are projective birational morphisms, the inclusions on the left are strict toroidal embeddings, $m_{\mathfrak{X}^{\circ}}^{-1}(\mathfrak{Z}_{\mathfrak{X}})$ is contained in the toroidal boundary $\mathfrak{X}^{\circ}\setminus U_{\mathfrak{X}^{\circ}}$, and $\Phi^{\circ}$ is a projective, surjective toroidal morphism of toroidal embeddings.
Step 2. Now we include the closed subset $\mathfrak{Z}_{\mathfrak{B}}$ into the toroidal boundary by an additional blowing up on $\mathfrak{B}^{\circ}$. Consider the closed subset $\mathfrak{Z}^{\circ}\subset\mathfrak{B}^{\circ}$ that is the union of the inverse image of $\mathfrak{Z}_{\mathfrak{B}}$ in $\mathfrak{B}^{\circ}$ and the toroidal boundary $\mathfrak{B}^{\circ}\setminus U_{\mathfrak{B}^{\circ}}$. Let $m_{\mathfrak{B}^{\prime}}\colon\mathfrak{B}^{\prime}\to\mathfrak{B}^{\circ}$ be a log resolution of $(\mathfrak{B}^{\circ},\mathfrak{Z}^{\circ})$ such that $\mathfrak{B}^{\prime}$ is smooth and the inverse image of $\mathfrak{Z}^{\circ}$ in $\mathfrak{B}^{\prime}$ is a simple normal crossing divisor. Set $U_{\mathfrak{B}^{\prime}}$ as the open complement $\mathfrak{B}^{\prime}\setminus m_{\mathfrak{B}^{\prime}}^{-1}(\mathfrak{Z}^{\circ})$. Then $(U_{\mathfrak{B}^{\prime}}\subset\mathfrak{B}^{\prime})$ is a strict toroidal embedding and
@@
where $m_{\mathfrak{X}^{\prime}}$ and $m_{\mathfrak{B}^{\prime}}$ are projective birational morphisms, $(U_{\mathfrak{X}^{\prime}}\subset\mathfrak{X}^{\prime})$ is also a strict toroidal embedding, the open subset $U_{\mathfrak{X}^{\prime}}\subset\mathfrak{X}^{\prime}$ is the intersection of the open subsets $(\Phi^{\prime})^{-1}(U_{\mathfrak{B}^{\prime}})$ and $m_{\mathfrak{X}^{\prime}}^{-1}(U_{\mathfrak{X}^{\circ}})$, and $\Phi^{\prime}$ is a surjective, projective toroidal morphism.
@@
$m_{\mathfrak{X}^{\circ}}\circ m_{\mathfrak{X}^{\prime}}\,\,\text{ and }\,\,m_{\mathfrak{B}^{\circ}}\circ m_{\mathfrak{B}^{\prime}}.$
@@
  • $m_{\mathfrak{B}}$ is an isomorphism over $\mathfrak{B}\setminus V_{\mathfrak{B}}$, and
  • every fibre of $\Phi$ over $\mathfrak{B}\setminus V_{\mathfrak{B}}$ is geometrically integral.
It is evident that $U_{\mathfrak{B}}\coloneqq\mathfrak{B}\setminus V_{\mathfrak{B}}$ is a non-empty open subset of $\mathfrak{B}$. Moreover, by construction, $U_{\mathfrak{B}}$ is contained in $\mathfrak{B}\setminus\mathfrak{Z}_{\mathfrak{B}}$ as the closed subset $\mathfrak{Z}_{\mathfrak{B}}$ is contained in $V_{\mathfrak{B}}$. Then the base change functoriality for curves in (ii) also follows from Theorem 3.11. ∎

### v3

@@ 3.8. Toroidal varieties and logarithmic geometry
As there are no other applications of logarithmic geometry in this paper except in the proof of Theorem 3.11, we refer the readers to [20, 30, 31, 39] for details about logarithmic geometry. For more details regarding the relation between toroidal geometry and logarithmic geometry, see [40].
All the log-structures in this paper are defined on Zariski sites. A log-scheme $(X,\mathcal{M})$ consists of a scheme $X$ and a log-structure $\alpha\colon\mathcal{M}\to\mathcal{O}_{X}$; see [31, §1]. A log-variety is a log-scheme whose underlying scheme is a variety. To avoid confusion, when saying logarithmically smooth log-morphisms, we mean logarithmically smooth morphisms in the category of log-schemes defined in the sense of [30, 31, 39]. However, when saying a log smooth morphism, we mean a log smooth morphism from a couple as in §2.8. The applications of these terminologies should be clear from the context.
Let $X$ be a variety and $\iota\colon U\hookrightarrow X$ an inclusion of an open subset. Denote by $\mathcal{M}_{U/X}$ the sheaf of monoids $\mathcal{O}_{X}\cap\iota_{*}\mathcal{O}_{X}^{*}$. We call $\mathcal{M}_{U/X}$ the compactifying log-structure associated to the embedding $(U\subset X)$; cf. [39, §III.1.6]. The log-variety $(X,\mathcal{M}_{U/X})$ is logarithmically regular (or log-regular for short) if and only if $(U\subset X)$ is a strict toroidal embedding; cf. [31, (1.7)] and [40].
Let $(U_{X}\subset X),(U_{Y}\subset Y)$ be strict toroidal embeddings. Let $(U_{X}\subset X)\to(U_{Y}\subset Y)$ be a morphism of embeddings, that is, a morphism $f\colon X\to Y$ of varieties such that $f(U_{X})\subseteq U_{Y}$. Then $(U_{X}\subset X)\to(U_{Y}\subset Y)$ is a dominant toroidal morphism if and only if the induced log-morphism $(X,\mathcal{M}_{U_{X}/X})\to(Y,\mathcal{M}_{U_{Y}/Y})$ of log-varieties is logarithmically smooth in the category of log-schemes; see [17, 40].
@@ 3.9. Relative toroidalisation
First we recall a theorem proved by Abramovich and Karu in [2], which says that any projective, surjective morphism of varieties can be made logarithmically smooth after birational modifications. This result is generalised to varieties over non-closed fields of characteristic zero in [1].
Theorem 3.10 ([1, 2]).
Let $f\colon X\to B$ be a projective, surjective morphism of varieties. Let $Z\subset X$ be a proper closed subset. Then there exists a commutative diagram
@@
We note that the method proving this theorem in [1, 2] is non-canonical and even a smooth generic fibre of $X\to B$ can be modified by it. In [3], it is shown that there is a relatively canonical toroidalisation method that keeps the toroidal locus of $f$ unchanged; for more details, we refer the readers to [3, §1]. In the next result, we show that a normalised base change of a dominant toroidal morphism is also toroidal. For more details and generalisation of this result to log-structures on small étale sites, we refer the readers to [40].
@@
$f\colon(U_{Y}\subset Y)\to(U_{Z}\subset Z)~~\mbox{and}~~g\colon(U_{X}\subset X)\to(U_{Z}\subset Z)$
@@
$f\colon(Y,\mathcal{M}_{U_{Y}/Y})\to(Z,\mathcal{M}_{U_{Z}/Z})~~\mbox{and}~~g\colon(X,\mathcal{M}_{U_{X}/X})\to(Z,\mathcal{M}_{U_{Z}/Z})$
@@
$\mathbf{\Omega}\coloneqq(\Omega,\mathcal{M}_{\Omega})~~\mbox{and}~~\mathbf{\Omega}^{\operatorname{fs}}\coloneqq(\Omega^{\operatorname{fs}},\mathcal{M}_{\Omega^{\operatorname{fs}}})$
the fibre product of $(Y,\mathcal{M}_{U_{Y}/Y})$ and $(X,\mathcal{M}_{U_{X}/X})$ along $(Z,\mathcal{M}_{U_{Z}/Z})$ in the category of log-schemes and fs log-schemes respectively. By [39, III.2.1.2], the underlying scheme $\Omega$ is isomorphic to $Y\times_{Z}X$, and $\mathcal{M}_{\Omega}$ is a coherent log-structure on $\Omega$.
By [20, Remark 12.2.36 (ii)], the morphism of schemes $\Omega^{\operatorname{fs}}\to\Omega$ is a finite surjective morphism. By [20, Proposition 12.3.34] and the proof of [20, Proposition 12.2.35],
@@
is log-étale. As logarithmic smoothness is stable under composition and base change (see [20, Proposition 12.3.24]), the induced log-morphism of log-schemes
@@
is logarithmically smooth. Then by [31, (8.2)], the log-scheme $\mathbf{\Omega}^{\operatorname{fs}}$ is log-regular. By [20, Corollary 12.5.29], $\Omega^{\operatorname{fs}}$ is a normal scheme, hence $\Omega^{\operatorname{fs}}$ is a disjoint union of normal varieties. Thus, every irreducible component of $\Omega^{\operatorname{fs}}$ is a strict toroidal variety.
Let $U$ be the fibre product $U_{Y}\times_{U_{Z}}U_{X}$. By [40, Lemma 6.2], $U$ is nonempty. Let $(U,\mathcal{M}_{U})$ be the log-scheme endowed with trivial log-structure. It is evident that $\mathcal{M}_{U}$ is the restriction of $\mathcal{M}_{\Omega}$ to $U$. Denote by $\Omega^{\operatorname{fs},*}$ the maximal open subset of $\Omega^{\operatorname{fs}}$ on which $\mathcal{M}_{\Omega^{\operatorname{fs}}}$ is trivial. Then by the construction in [20, Proposition 12.2.35], $U$ is also an open subscheme of $\Omega^{\operatorname{fs}}$ satisfying $\mathcal{M}_{U}=\mathcal{M}_{\Omega^{\operatorname{fs}}}|_{U}$, whence $U\subseteq\Omega^{\operatorname{fs},*}$. Moreover, by [20, Corollary 12.3.27], the scheme $U$ is nonsingular in the usual sense.
Step 2. Let $W^{\prime}$ be an irreducible component of $Y\times_{Z}X$ that dominates $Y$. By [40, Lemma 6.2], the inverse image of $U_{X}$ in $W^{\prime}$ is a dense open subset of $W^{\prime}$. Let $W$ be the normalisation of $W^{\prime}$, and let $p\colon W\to Y$ and $q\colon W\to X$ be the projections. Then $p^{-1}(U_{Y})\cap q^{-1}(U_{X})$ is a nonempty dense open subset of $W$. Since $U$ is a nonsingular scheme, there is a unique irreducible component $U_{W^{\prime}}$ of $U$ contained in $W^{\prime}$. Then the normalisation $W\to W^{\prime}$ is an isomorphism over $U_{W^{\prime}}$. Denote by $U_{W}$ the inverse image of $U_{W^{\prime}}$ in $W$. Then $U_{W}=p^{-1}(U_{Y})\cap q^{-1}(U_{X})$. Let $W^{\prime\prime}$ be the irreducible component of $\Omega^{\operatorname{fs}}$ that contains $U_{W^{\prime}}$. As $U_{W^{\prime}}$ embeds into both $W^{\prime}$ and $W^{\prime\prime}$, the finite morphism $W^{\prime\prime}\to W^{\prime}$ is also birational. Since $\Omega^{\operatorname{fs}}$ is a normal scheme, we see that $W^{\prime\prime}\cong W$.
Denote by $W^{*}\subseteq W$ the restriction of $\Omega^{\operatorname{fs},*}$ to $W$. By [20, Remark 12.2.8], $W^{*}$ is mapped into $U_{Y},U_{X}$ via $p\colon W\to Y,q\colon W\to X$ respectively. Then
@@
    [Base change functoriality] There is a nonempty open subset $U_{\mathfrak{B}}\subset\mathfrak{B}\setminus\mathfrak{Z}_{\mathfrak{B}}$ satisfying the following property. Let $i\colon C\to\mathfrak{B}$ be a morphism from a smooth curve. Assume that the image $i(C)$ is not entirely contained in the closed subset $\mathfrak{B}\setminus U_{\mathfrak{B}}$. Then $\mathfrak{X}\times_{\mathfrak{B}}C$ admits a unique irreducible component dominating $C$. Denote by $Y$ the reduction of this irreducible component of $\mathfrak{X}\times_{\mathfrak{B}}C$ with projection morphism $f\colon Y\to C$. Consider the diagram arising from normalised base change as follows:
@@
    where $C^{\prime}$ (respectively, $Y^{\prime}$) is the normalisation of the main component of $\mathfrak{B}^{\prime}\times_{\mathfrak{B}}C$ (respectively, of $\mathfrak{X}^{\prime}\times_{\mathfrak{B}^{\prime}}C^{\prime}$), and $U_{C^{\prime}}$ (respectively, $U_{Y^{\prime}}$) is the inverse image of $U_{\mathfrak{B}^{\prime}}$ in $C^{\prime}$ (respectively, of $U_{\mathfrak{X}^{\prime}}$ in $Y^{\prime}$). Then $U_{C^{\prime}},U_{Y^{\prime}}$ are nonempty open subsets of $C^{\prime},Y^{\prime}$ respectively, $f^{\prime}$ is also a toroidal morphism of strict toroidal embeddings, and all the properties in (i) are satisfied. In particular, $m_{C}$ and $m_{Y}$ are projective birational morphisms, and the inverse image of $\mathfrak{Z}_{\mathfrak{B}}$ in $C^{\prime}$ (respectively, of $\mathfrak{Z}_{\mathfrak{X}}$ in $Y^{\prime}$) is contained in $C^{\prime}\setminus U_{C^{\prime}}$ (respectively, in $Y^{\prime}\setminus U_{Y^{\prime}}$).
@@
Step 1. To include the closed subset $\mathfrak{Z}_{\mathfrak{X}}$ into the toroidal structure, we first modify $\Phi\colon\mathfrak{X}\to\mathfrak{B}$ by Theorem 3.10 as in the following commutative diagram
@@
where $m_{\mathfrak{X}^{\circ}}$ and $m_{\mathfrak{B}^{\circ}}$ are projective birational morphisms, the inclusions on the left are strict toroidal embeddings, $m_{\mathfrak{X}^{\circ}}^{-1}(\mathfrak{Z}_{\mathfrak{X}})$ is contained in the toroidal boundary $\mathfrak{X}^{\circ}\setminus U_{\mathfrak{X}^{\circ}}$, and $\Phi^{\circ}$ is a projective, surjective toroidal morphism of strict toroidal embeddings.
Step 2. Now we include the closed subset $\mathfrak{Z}_{\mathfrak{B}}$ into the toroidal boundary by an additional modification on $\mathfrak{B}^{\circ}$. Consider the closed subset $\mathfrak{Z}^{\circ}\subset\mathfrak{B}^{\circ}$ that is the union of the inverse image of $\mathfrak{Z}_{\mathfrak{B}}$ in $\mathfrak{B}^{\circ}$ and the toroidal boundary $\mathfrak{B}^{\circ}\setminus U_{\mathfrak{B}^{\circ}}$. Let $m_{\mathfrak{B}^{\prime}}\colon\mathfrak{B}^{\prime}\to\mathfrak{B}^{\circ}$ be a log resolution of $(\mathfrak{B}^{\circ},\mathfrak{Z}^{\circ})$ such that $\mathfrak{B}^{\prime}$ is smooth and the inverse image of $\mathfrak{Z}^{\circ}$ in $\mathfrak{B}^{\prime}$ is a simple normal crossing divisor. Set $U_{\mathfrak{B}^{\prime}}$ as the open complement $\mathfrak{B}^{\prime}\setminus m_{\mathfrak{B}^{\prime}}^{-1}(\mathfrak{Z}^{\circ})$. Then $(U_{\mathfrak{B}^{\prime}}\subset\mathfrak{B}^{\prime})$ is a strict toroidal embedding and
@@
where $\mathfrak{X}^{\prime}$ is the normalisation of the main component of $\mathfrak{B}^{\prime}\times_{\mathfrak{B}^{\circ}}\mathfrak{X}^{\circ}$, $m_{\mathfrak{X}^{\prime}}$ and $m_{\mathfrak{B}^{\prime}}$ are projective birational morphisms, $(U_{\mathfrak{X}^{\prime}}\subset\mathfrak{X}^{\prime})$ is also a strict toroidal embedding, the open subset $U_{\mathfrak{X}^{\prime}}\subset\mathfrak{X}^{\prime}$ is the intersection of the open subsets $(\Phi^{\prime})^{-1}(U_{\mathfrak{B}^{\prime}})$ and $m_{\mathfrak{X}^{\prime}}^{-1}(U_{\mathfrak{X}^{\circ}})$, and $\Phi^{\prime}$ is a surjective, projective toroidal morphism.
@@
$m_{\mathfrak{X}^{\circ}}\circ m_{\mathfrak{X}^{\prime}}~~\mbox{and}~~m_{\mathfrak{B}^{\circ}}\circ m_{\mathfrak{B}^{\prime}}.$
@@
  • $m_{\mathfrak{B}}$ is an isomorphism over $\mathfrak{B}\setminus V_{\mathfrak{B}}$,
  • every closed fibre of $\Phi$ over $\mathfrak{B}\setminus V_{\mathfrak{B}}$ is integral, and
  • for every closed point $b\in\mathfrak{B}\setminus V_{\mathfrak{B}}$, let $b^{\prime}$ be its inverse image in $\mathfrak{B}^{\prime}$, then $m_{\mathfrak{X}}$ induces a birational morphism $\mathfrak{X}^{\prime}_{b^{\prime}}\to\mathfrak{X}_{b}$ on the fibres (in particular, $\mathfrak{X}^{\prime}_{b^{\prime}}$ is irreducible).
It is evident that $U_{\mathfrak{B}}\coloneqq\mathfrak{B}\setminus V_{\mathfrak{B}}$ is a nonempty open subset of $\mathfrak{B}$. Moreover, by construction, $U_{\mathfrak{B}}$ is contained in $\mathfrak{B}\setminus\mathfrak{Z}_{\mathfrak{B}}$ as the closed subset $\mathfrak{Z}_{\mathfrak{B}}$ is contained in $V_{\mathfrak{B}}$.
Let $i\colon C\to\mathfrak{B}$ be a morphism from a smooth curve such that the image $i(C)$ is not entirely contained in $V_{\mathfrak{B}}$. As $m_{\mathfrak{B}}$ is an isomorphism over $U_{\mathfrak{B}}$, $\mathfrak{B}^{\prime}\times_{\mathfrak{B}}C$ admits a unique irreducible component that is birational to $C$. Denote by $C^{\prime}$ the normalisation of this irreducible component, then the induced morphism $C^{\prime}\to C$ is an isomorphism of nonsingular curves. Let $U_{C^{\prime}}$ be the inverse image of $U_{\mathfrak{B}^{\prime}}$ in $C^{\prime}$, which is nonempty as $V_{\mathfrak{B}}$ contains the image of $\mathfrak{B}^{\prime}\setminus U_{\mathfrak{B}^{\prime}}$. It is evident that $(U_{C^{\prime}}\subset C^{\prime})$ is a strict toroidal embedding. Note that $U_{C^{\prime}}$ may be equal to the whole $C^{\prime}$.
By the construction of $U_{\mathfrak{B}}$, a general closed fibre of $\mathfrak{X}\times_{\mathfrak{B}}C\to C$ is irreducible, hence there is a unique irreducible component $X$ of $\mathfrak{X}\times_{\mathfrak{B}}C$ that dominates $C$. For the same reason, $\mathfrak{X}^{\prime}\times_{\mathfrak{B}^{\prime}}C^{\prime}$ admits a unique irreducible component $X^{\prime}$ that dominates $C^{\prime}$. Moreover, $m_{\mathfrak{X}}$ induces a birational morphism $X^{\prime}\to X$. By [22, Chap. I, (5.1.5)], there is an induced birational morphism $X^{\prime}_{\operatorname{red}}\to X_{\operatorname{red}}$. Let $Y\coloneqq X_{\operatorname{red}}$, and let $Y^{\prime}$ be the normalisation of $X^{\prime}_{\operatorname{red}}$. Then it is evident that the induced morphism $m_{Y}\colon Y^{\prime}\to Y$ is projective and birational.
Let $f^{\prime}$ be the morphism $Y^{\prime}\to C^{\prime}$. Let $U_{Y^{\prime}}$ be the inverse image of $U_{\mathfrak{X}^{\prime}}$ in $Y^{\prime}$. Note that $U_{Y^{\prime}}\cap(f^{\prime})^{-1}(U_{C^{\prime}})$ is equal to $U_{Y^{\prime}}$ as the inverse image of $U_{\mathfrak{B}^{\prime}}$ in $\mathfrak{X^{\prime}}$ contains $U_{\mathfrak{X}^{\prime}}$; see Lemma 3.7. Applying Theorem 3.11 to the toroidal morphism $\Phi^{\prime}$ and the morphism of embeddings $(U_{C^{\prime}}\subset C^{\prime})\to(U_{\mathfrak{B}^{\prime}}\subset\mathfrak{B}^{\prime})$, we see that $U_{Y^{\prime}}$ is nonempty, $(U_{Y^{\prime}}\subset Y^{\prime})$ is also a strict toroidal embedding, and the induced morphism of embeddings $f^{\prime}\colon(U_{Y^{\prime}}\subset Y^{\prime})\to(U_{C^{\prime}}\subset C^{\prime})$ is a toroidal morphism of strict toroidal embeddings. This concludes (ii). ∎